The atomic number of Ann element x is 14 and it's relative atomic mass is 28. How many neutrons does an atom x contains? Hence write it's electronic configuration


1
Expert's answer
2021-10-26T05:37:34-0400

14 neutrons - relative atomic mass = number of protons (given by the atomic number) + number of neutrons.

So for RAM = 28 = NP + NN = 14 + x => x = 14

Element with atomic number 14 is Si. Atom Si contain 28 - 14 = 14 neutrons. Its electronic configuration is 1s22s22p63s23s2.